OpenStudy (anonymous):

what is value of x? 2x+3/5x+3=2/5 a)x=1 b)x=3 c) no real numbers d) all real numbers

OpenStudy (anonymous):

\[2x+\frac{3}{5}x+3=\frac{2}{5}\] take out x on LHS as common factor (2+0.6)x+3=2/5 2.6x+3=0.4 2.6x=-2.6 x=-1 (not 1)

OpenStudy (anonymous):

Thanks for explaining that!